Calculate the molarity of 20.5 g sodium chloride (NaCl) compound with molar mass of 

58.44g/mole placed in 55.0 mL of total solution. 



1
Expert's answer
2021-03-16T08:25:33-0400

Moles of NaCl = "\frac{20.5}{58.44}=0.35moles"


"55.0mL = \frac{55}{1000}=0.055 L"


Hence


"Molarity = \frac{0.35}{0.055}=6.36M"